What is the maximum number of occupants you must provide a means of egress for a building in an uncovered mall with 500,000 sq ft of leasable space? assume that the build out for the space took up 45% of the leasable space.

60

61

59

50
Answer #1

We know that occupant load factor OLF is calculated by the factor

OLF = (0.00007) (GLA) + 25

where GLA is the gross leasable area

OLF = occupant load factor = .00007GLA + 25

So we have OLF = (.00007*500000) + 25

= 35 + 25 = 60

So answer is Option A = 60

===========================================================================================

Please vote thumbs up to support us.